# Propagation of Chaos for Mean-Field Langevin Dynamics and its Application to Model Ensemble

### v1

@@ 5.2 LoRA for finetuning language models
In recent years, efforts to improve predictive capabilities and computational efficiency in machine learning have revived interest in techniques such as ensembling (Ganaie et al., 2022; Mohammed & Mohammed, 2023) and model merging (Yang et al., 2024; Goddard et al., 2024). Ensemble methods improve predictive performance by combining the outputs of multiple models during inference (Hansen & Salamon, 1990; Dietterich, 2000). Although several fusion variants exist (Kim et al., 2003; Soares et al., 2004), Cheng et al. (2018) shows that simple average voting (Breiman, 1996) does not perform significantly worse while still being highly efficient, with uses in several deep learning applications (Cheng et al., 2018; Vazquez-Romero & Gallardo-Antolin, 2020).
In contrast, model merging consolidates multiple models into a single one by combining individual parameters, showing success particularly in the optimization of LLMs (Ilharco et al., 2022; Jin et al., 2023; Davari & Belilovsky, 2024). An approach to merging models is to simply average the weights across multiple models (Utans, 1996). Taking the average of weights along a single optimization trajectory has been demonstrated to achieve better generalization (Izmailov et al., 2018; Frankle et al., 2020). Moreover, interpolating any two random weights from models that lie in the same loss basins could produce even more optimal solutions that are closer to the centre of the basin (Neyshabur et al., 2020). These works then form the foundation of model soups in Wortsman et al. (2022) which refers to averaging the weights of independently fine-tuned models. Similarly, Gauthier-Caron et al. (2024) showed that basic weight averaging methods can perform competitively if constituent models are similar, despite the emergence of novel LLM merging strategies (Ramé et al., 2023; Chronopoulou et al., 2023; Yu et al., 2024).
Despite the widespread use of model merging in the current research landscape, theoretical results concerning the merging of fully trained neural networks are limited. For models trained with stochastic gradient descent, averaging the model weights during different iterations of a single run improves stability bounds (Hardt et al., 2016) and variance (Jain et al., 2018) under convex assumptions. Stability bounds in the non-convex settings are then addressed by Wang et al. (2024). Ortiz-Jimenez et al. (2023) studied weight interpolation techniques for task arithmetic in vision-language models, demonstrating that linearized models under the neural tangent kernel regime can outperform non-linear counterparts.

### v2

@@ Acknowledgements
In recent years, efforts to improve predictive capabilities and computational efficiency in machine learning have revived interest in techniques such as ensembling (Ganaie et al.,, 2022; Mohammed and Mohammed,, 2023) and model merging (Yang et al.,, 2024; Goddard et al.,, 2024). Ensemble methods improve predictive performance by combining the outputs of multiple models during inference (Hansen and Salamon,, 1990; Dietterich,, 2000). Although several fusion variants exist (Kim et al.,, 2003; Soares et al.,, 2004), Cheng et al., (2018) shows that simple average voting (Breiman,, 1996) does not perform significantly worse while still being highly efficient, with uses in several deep learning applications (Cheng et al.,, 2018; Vazquez-Romero and Gallardo-Antolin,, 2020).
In contrast, model merging consolidates multiple models into a single one by combining individual parameters, showing success particularly in the optimization of LLMs (Ilharco et al.,, 2022; Jin et al.,, 2023; Davari and Belilovsky,, 2024). An approach to merging models is to simply average the weights across multiple models (Utans,, 1996). Taking the average of weights along a single optimization trajectory has been demonstrated to achieve better generalization (Izmailov et al.,, 2018; Frankle et al.,, 2020). Moreover, interpolating any two random weights from models that lie in the same loss basins could produce even more optimal solutions that are closer to the centre of the basin (Neyshabur et al.,, 2020). These works then form the foundation of model soups in Wortsman et al., (2022) which refers to averaging the weights of independently fine-tuned models. Similarly, Gauthier-Caron et al., (2024) showed that basic weight averaging methods can perform competitively if constituent models are similar, despite the emergence of novel LLM merging strategies (Ramé et al.,, 2023; Chronopoulou et al.,, 2023; Yu et al.,, 2024).
Despite the widespread use of model merging in the current research landscape, theoretical results concerning the merging of fully trained neural networks are limited. For models trained with stochastic gradient descent, averaging the model weights during different iterations of a single run improves stability bounds (Hardt et al.,, 2016) and variance (Jain et al.,, 2018) under convex assumptions. Stability bounds in the non-convex settings are then addressed by Wang et al., (2024). Ortiz-Jimenez et al., (2023) studied weight interpolation techniques for task arithmetic in vision-language models, demonstrating that linearized models under the neural tangent kernel regime can outperform non-linear counterparts.